Find det(C) if

C = λ λ + 1

λ λ − 1


1
Expert's answer
2021-05-14T10:57:43-0400

If

"C = \left( {\begin{matrix}
\lambda &{\lambda + 1}\\
\lambda &{\lambda - 1}
\end{matrix}} \right)"

Then

"\det (C) = \lambda \left( {\lambda - 1} \right) - \lambda \left( {\lambda + 1} \right) = {\lambda ^2} - \lambda - {\lambda ^2} - \lambda = - 2\lambda"

Answer: "- 2\lambda"